How much potential energy does a car gain if a crain lifts the car with a mass of 1,500 kg and 20 meters straight up?

1
Expert's answer
2022-01-05T14:30:02-0500

The potential energy is given as follows:


"E = mgh"

where "m = 1500kg, h=20m, g = 9.8m\/s^2". Thus, obtain:


"E = 1500\cdot 20\cdot 9.8 = 2.94\times 10^5J"

Answer. "2.94\times 10^5J".